# Real-World Data and Calibrated Simulation Suite for Offline Training of Reinforcement Learning Agents to Optimize Energy and Emission in Buildings for Environmental Sustainability

### v1

@@ 4 The Dataset Structure
Both the real-world data and simulated data are given in the same format. Following the RL paradigm, data is provided as a series of observations, actions, and rewards. In the case of the real-world data, this comes in the form of static historical episodes, where the actions follow the baseline policy in the building, and in the case of the simulator, as a proper interactive RL environment where actions can be taken in real time.
To make the task as realistic as possible, we formatted the data to closely resemble the real-world building API, so that a user can mimic interacting with the building. All of our data is formatted to be compliant with the popular open source Google Digital Buildings Ontology (DBO). The agent communicates with the building using the Protobuf open source serialization format(Google, ). The agent can send information requests to the building, asking for structural information, such as the number of devices, and telemetry information, such as the value of a particular sensor, and the building sends back a response, containing the requested information. The agent can also request that a setpoint be changed to a new value, and the building will respond if the change was successful.
Following the RL paradigm, the data in our dataset falls under the following categories:
@@
    Environment Data or each building environment, the dataset contains information on all HVAC zones and HVAC devices. For zones this includes the name and size of each zone, as well has how many devices are contained within it. For devices, this includes the zone the device is associated with, as well as every device sensor and setpoint.
@@
SB1 93,858 2 170
SB2 62,613 1 152
SB3 118,086 3 152
The dataset currently consists of six years of data from three buildings. The details are in Table 1. For more details regarding the format of the data, including definitions and examples of each type of proto, see appendix B.
Data Visualization We also present a data visualization module, compatible both for viewing the real-world historical data, as well as visualizing the state of the simulator, as shown in Figure 1. Given an observation of a building environment, our visualization module renders a two dimensional heat-map view of the building. This greatly aids in understanding the data, and is invaluable in understanding how a particular policy is behaving.
5 Simulator Design Considerations
Every building is unique, due to its physical layout, equipment, and location. Fully customizing a high fidelity simulation to a specific target building requires nearly exhaustive knowledge of the building structure, materials, location, etc., some of which are unknowable, especially for legacy office buildings. This requires manual “guesstimation”, which can erode the accuracy promised by high-fidelity simulation. In general, the configuration time required for high-fidelity simulations limits their utility for deploying RL-based optimization to many buildings. High-fidelity simulations also are affected by computational demand and long execution times.
Alternatively, we propose a fast, low-to-medium-fidelity simulation model that was useful in addressing various design decisions, such as the reward function, the modeling of different algorithms. and for end-to-end testing. The simulation is built on a 2D finite-difference (FD) grid that models thermal diffusion, and a simplified HVAC model that generates or removes heat on special “diffuser” control volumes (CV) in the FD grid. For more details on design considerations, see Appendix C.
While the uncalibrated simulator is of low-to-medium fidelity, the key additional factor is data. We collect recorded observations from the target building under baseline control, and use that data to calibrate the simulator, by adjusting the simulator’s physical parameters to minimize difference between real and simulated data. We believe this approach hits the sweet spot in this tradeoff, enabling scalability, while maintaining a high enough level of fidelity to train an improved policy.
Figure 2: Simulation Fidelity vs. Execution Speed. The ideal operating point for training RL agents for energy and emission efficiency is a tradeoff between fidelity, depicted as 1 minus a normalized error $\epsilon$ between simulation and real, and execution speed, as measured by the number of training steps per second. Additional consideration also includes the time to configure a custom simulator for the target building. While many approaches tend to favor high-fidelity over execution, speed, our approach argues a low-to-medium fidelity that has a medium-to-high speed is most suitable for training an RL agent.
6 A Lightweight, Calibrated Simulation
Our goal is to develop a method for applying RL at scale to commercial buildings. To this end, we put forth the following requirements for this to be feasible: We must have an easily customizable simulated environment to train the agent, with high enough fidelity to train an improved control agent. To meet these desiderata, we designed a light weight simulator based on finite differences approximation of heat exchange, building upon earlier work (Goldfeder & Sipple, 2023). We proposed a simple automated procedure to go from building floor plans to a custom simulator in a short time, and we designed a calibration and evaluation pipeline, to use data to fine tune the simulation to better match the real world. What follows is a description of our implementation.
@@ 6 A Lightweight, Calibrated Simulation
We selected water supply temperature $\hat{T}_{b}$ and the air handler supply temperature $\hat{T}_{s}$ as agent actions because they affect the balance of electricity and natural gas consumption, they affect multiple device interactions, and they affect occupant comfort. Greater efficiencies can be achieved with these setpoints by choosing the ideal times and values to warm up and cool down the building in the workday mornings and evenings. Further tradeoffs include balancing the thermal load between hot water heating with natural gas and supply air heating with electricity using the air conditioner or heat pump units.
@@
Finite Differences Approximation The diffusion of thermal energy in time and space of the building can be approximated using the method of Finite Differences (FD)(Sparrow, 1993; Lomax et al., 2002), and applying an energy balance. This method divides each floor of the building into a grid of three-dimensional control volumes and applies thermal diffusion equations to estimate the temperature of each control volume. By assuming each floor is adiabatically isolated, (i.e., no heat is transferred between floors), we can simplify the three-spatial dimensions into a spatial two-dimensional heat transfer problem. Each control volume is a narrow volume bounded horizontally, parameterized by $\Delta x^{2}$, and vertically by the height of the floor. The energy balance, shown below, is applied to each discrete control volume in the FD grid, and consists of the following components: (a) the thermal exchange across each face of the four participating faces control volume via conduction or convection $Q_{1}$, $Q_{2}$, $Q_{3}$, $Q_{4}$, (b) the change in internal energy over time in the control volume $Mc\frac{\Delta T}{\Delta t}$, and (c) an external energy source that enables applying local thermal energy from the HVAC model only for those control volumes that include an airflow diffuser, $Q_{ext}$. The equation is $Q_{ext}+Q_{1}+Q_{2}+Q_{3}+Q_{4}=Mc\frac{\Delta T}{\Delta t}$, where $M$ is the mass and $c$ is the heat capacity of the control volume, $\Delta T$ is the temperature change from the prior timestep and $\Delta t$ is the timestep interval.
@@
Figure 3: Thermal model for simulation. A building consists of conditioned zones, where the mean temperature of the zone $T_{z}$ should be within upper and lower setpoints, $\hat{T}_{z,max}$ and $\hat{T}_{z,min}$. Thermal power for heating or cooling the room is supplied to each zone, $\dot{Q}_{s}$, and recirculated from the zone, $\dot{Q}_{r}$ from the HVAC system, with additional thermal exchange $\dot{Q}_{z}$ from walls, doors, etc. The Air Handler supplies the building with air at supply air temperature setpoint $\hat{T}_{s}$ drawing fresh air, $\dot{m}_{amb}$, at ambient temperatures, $T_{amb}$, and returning exhaust air $\dot{m}_{exhaust}$ at temperature $T_{exhaust}$ to the outside using intake and exhaust fans, $\dot{W}_{a,in}$ and $\dot{W}_{a,out}$. A fraction of the return air can be recirculated, $\dot{m}_{recirc}$. Central air conditioning is achieved with a chiller and pump that joins a refrigeration cycle to the supply air, consuming electrical energy for the AC compressor $\dot{W}_{c}$ and coolant circulation, $\dot{W}_{c,p}$. The hot water cycle consists of a boiler that maintains the supply water temperature at $T_{b}$ heated by natural gas power $\dot{Q}_{b}$, and a pump that circulates hot water through the building, with electrical power $\dot{W}_{b,p}$. Supply air is delivered to the zones through Variable Air Volume (VAV) devices.
@@
The thermal exchange in (a) is calculated using Fourier’s law of steady conduction in the interior control volumes (walls and interior air), parameterized by the conductivity of the volume, and the exchange across the exterior faces of control volumes are calculated using the forced convection equation, parameterized by the convection coefficient, which approximates winds and currents surrounding the building. The change in internal energy (b) is parameterized by the density, and heat capacity of the control volume. Finally, the thermal energy associated with the VAV (c) is equally distributed to all associated control volumes that have a diffuser. Thermal diffusion within the building is mainly accomplished via forced or natural convection currents, which can be notoriously difficult to estimate accurately. We note that heat transfer using air circulation is effectively the exchange of air mass between control volumes, which we approximate by a randomized shuffling of air within thermal zones, parameterized by a shuffle probability and radius. For more details on this approximation and associated equations, see Appendix D.
@@
Simulator Configuration For RL to scale to many buildings, it is critical to be able to easily and rapidly configure the simulator to any arbitrary building. We designed a procedure that, given floorplans and HVAC layout information, enables generating a fully specified simulation very rapidly. For example, on SB1, consisting of two floors and 170 devices, a single technician was able to configure the simulator in under three hours. Details of this procedure are provided in Appendix E.
@@
Simulator Calibration and Evaluation In order to calibrate the simulator to the real world using data, we must have a metric with which to evaluate our simulator’s fidelity, and an optimization method to improve our simulator on this metric.
@@
$\bm{N}$-Step Evaluation We propose a novel evaluation procedure, based on $N$-step prediction. Each iteration of our simulator was designed to represent a five-minute interval, and our real-world data is also obtained in five-minute intervals. To evaluate the simulator, we take a chunk of real data, consisting of $N$ consecutive observations. We then initialize the simulator so that its initial state matches that of the starting observation, and run the simulator for $N$ steps, replaying the same HVAC policy as was used in the real world. We then calculate our simulation fidelity metric, which is the mean absolute error of the temperatures in each temperature sensor at each time step, averaged over time. More formally, we define the Temporal Spatial Mean Absolute Error (TS-MAE) of $Z$ zones over $N$ timesteps as:
@@
Where $T_{real,t,z}$ is the measured zone air temperature for zone $z$ at timestamp $t$, and $T_{sim,t,z}=\frac{1}{|C_{z}|}\sum^{C_{z}}_{c=1}T_{t,c}$ is the mean temperature of all control volumes $C_{z}$ in zone $z$ at time $t$.
Hyperparameter Calibration Once we defined our simulation fidelity metric, the TS-MAE, we can attempt to minimize this error, thus improving fidelity, by hyperparameter tuning several physical constants and other variables using black-box optimization methods. We chose the method outlined in Golovin et. al. (Golovin et al., 2017), which automatically chooses the most appropriate strategy from a variety of popular algorithms.
7 Simulator Calibration
We now provide a full end-to-end demonstration of our calibration procedure, and show that our simulator, when tuned and calibrated, is able to make useful real-world predictions, and can train an RL agent to produce an improved policy over the baseline.
Setup We calibrated the simulator using data from SB1, with two stories, a combined surface area of 93,858 square feet, and 170 HVAC devices. Using the configuration pipeline, we went from floor plan blueprints to a fully configured simulator for this building, a process that took a single technician less than three hours to complete.
Calibration Data To calibrate our simulator, we took real-world data from three days, from Monday July 10, 2023 12:00 AM PST, to Thursday July 13, 2023 12:00 AM PST. The first two days were used as a train set, and the third day as validation of the calibrated performance on unseen data, as can be seen in Table 2. All times are given in US Pacific, the local time of the real building.
Calibration Procedure We ran hyperparameter tuning for 4000 iterations, with the aim of optimizing the TS-MAE, as outlined in equation 1, over the train data. We reviewed the physical constants that yielded the lowest simulation error from calibration. Densities, heat capacities, and conductivities plausibly matched common interior and exterior building materials. However, the external convection coefficient was higher than under the weather conditions, and likely is compensating for the radiative losses and gains, which were not directly simulated. For details about the hyperparameter tuning procedure, including the parameters varied, the ranges given, and the values found that best minimized the calibration metric, see Appendix F.
Calibration Results In Table 2, we present the predictive results of our calibrated simulator, on $N$-step prediction, for the train scenario, where $N=576$, representing a two day predictive window, and the test scenario, where $N=288$, representing a one day window. We calculated the TS-MAE, as defined in equation 1. We show results for the hyperparameters that best fit the train set, as well as for an uncalibrated simulator as a baseline. At no point was the validation data ever provided to the tuning process. Note that the validation period is half the duration of the train period, so a lower error does not mean we are performing better than on the train data.
Table 2: Training and test data scenarios
Split Length Start End calibrated $\epsilon$ uncalibrated $\epsilon$
train 48 hrs 2023-07-10 12AM 2023-07-12 12AM 0.717 ${}^{\circ}C$ 1.971 ${}^{\circ}C$
val. 24 hrs 2023-07-12 12AM 2023-07-13 12AM 0.566 ${}^{\circ}C$ 1.618 ${}^{\circ}C$
As indicated in Table 2, our tuning procedure drifts only $0.56\circ C$ on average over a 24-hour period on the validation set.
Visualizing Temperature Drift Over Time Figure 5 illustrates temperature drift over time for the training scenario. At each time step, we calculate the spatial temperature for all sensors in both the real building and simulator, and present them as side-by-side boxplot distributions for comparison. Figure 5 shows the same for the validation scenario.
@@ 7 Simulator Calibration
Here we can see that our simulator temperature distribution maintains a minimal drift from the real world, although it does seem a bit less reactive to daily fluctuation patterns, which may be the result of the lack of a radiative heat transfer model.
Visualizing Spatial Errors Figure 6 illustrates the results of this predictive process over a 24-hour period, on the validation data. It displays a heatmap of the spatial temperature difference throughout the building, between the real world and simulator, after 24 hours of the simulator making predictions. The ring of blue around the building indicates that our simulator is too cold on the perimeter, which implies that the heat exchange with the outside is happening more rapidly than it would in the real world. The inside of the building, at least on the first floor, contains significant amounts of red, indicating that despite the simulator perimeter being cooler than the real world, the inside is warmer. This implies that our thermal exchange within the building is not as rapid as that of the real world. We suspect that this may be because our simulator does not have a radiative heat transfer model. Lastly, there is a large amount of white in this image, indicating that for the most part, even after 24 hours of making predictions on the validation data, our calibration process was successful and the fidelity remains high. For more visuals of spatial errors, see appendix G.
@@
Figure 6: Visualization of simulator drift after 24 hours, on the validation data. The image is a heat map representing the temperature difference between the simulator and the real world, with red indicating the simulator is hotter, blue indicating it is colder, and white indicating no difference. The zone with the max and min temperature difference are indicated by displaying above them the difference.
8 Demonstration Benchmarking Results
While we believe our benchmark will be useful for a variety of tasks, such as further use of the data to calibrate the simulator, in this section we highlight results on three important tasks that our suite is well suited to: training an RL agent on the simulator, training a time-series regression model to predict the real world data, and training an RL agent on the real data directly.

### v2

@@ 4 The Smart Buildings Dataset
The real-world data and interactive simulated data are given in the same format. Following the RL paradigm, data are provided as a series of observations, actions, and rewards. The real-world data are in the form of static historical episodes, where actions follow the baseline policy in the building. The simulator and PINN are interactive RL environments where actions can be taken in realtime. The data fall into the following categories:
@@
    Environment Data For each building environment, the dataset contains information on all zones and devices. This includes the name and size of each zone, and the zone, location, sensors, and setpoints of each device.
@@
SB1 93,858 2 173
SB2 62,613 1 144
SB3 118,086 3 281
SB4 50,852 6 128
SB5 5,500 2 11
SB6 5,120 1 6
SB7 76,000 5 148
SB8 74,631 5 119
SB9 90,650 5 136
SB10 86,150 4 231
SB11 58,217 4 101
The data is from 11 buildings across North America, spans 3 building management systems, and was collected over a 6-year period. This diversity makes it ideal for building environment models that scale. For information on the data format and how to access the benchmark, see Appendix C.
Data Visualization We also present a data visualization module, compatible both for viewing the real-world data, as well as visualizing the state of the simulator, as shown in Figure 1. Given an observation of a building environment, our visualization module renders a two-dimensional heatmap view of the building. This greatly aids in understanding the data and analyzing how a particular policy is behaving.
5 The Smart Buildings Simulator
Our goal is to develop a method for applying RL at scale to commercial buildings. To this end, we put forth the following requirements for this to be feasible: We must have a lightweight, easy to configure and calibrate simulated environment to train the agent, with high enough fidelity to train an improved control agent. To meet these desiderata, we designed a lightweight simulator based on finite differences approximation of heat exchange, building upon earlier work (Goldfeder & Sipple, 2023). We proposed a simple automated procedure to go from building floor plans to a custom simulator in a short time, and we designed a calibration and evaluation pipeline, to use data to fine tune the simulation to better match the real world. What follows is a description of our implementation. For details regarding design considerations, see Appendix D.
@@ 5 The Smart Buildings Simulator
We selected hot water supply temperature $\hat{T}_{b}$ and the air handler supply temperature $\hat{T}_{s}$ as agent actions because they affect the balance of electricity and natural gas consumption, they affect multiple device interactions, and they affect occupant comfort. Greater efficiencies can be achieved with these setpoints by choosing the ideal times and values to warm up and cool down the building in the workday mornings and evenings. Further tradeoffs include balancing the thermal load between hot water heating with natural gas and supply air heating with electricity using the AC or heat pump units.
@@
Finite Differences Approximation The diffusion of thermal energy in time and space of the building can be approximated using the method of Finite Differences (FD)(Sparrow, 1993; Lomax et al., 2002), and applying an energy balance. This method divides each floor of the building into a grid of three-dimensional control volumes (CVs) and applies thermal diffusion equations to estimate the temperature of each CV. By assuming each floor is adiabatically isolated, we can simplify the three spatial dimensions into a spatial two-dimensional heat transfer problem. Each CV is a narrow volume bounded horizontally, parameterized by $\Delta x^{2}$, and vertically by the floor height. The energy balance, shown below, is applied to each discrete CV in the FD grid and consists of the following components: (a) the thermal exchange across each face of the four participating neighbor CVs via conduction or convection $Q_{1}$, $Q_{2}$, $Q_{3}$, $Q_{4}$, (b) the change in internal energy over time in the CV $Mc\frac{\Delta T}{\Delta t}$, and (c) an external energy source that enables applying local thermal energy from the HVAC model only for those CVs that include an airflow diffuser, $Q_{ext}$. The equation is $Q_{ext}+Q_{1}+Q_{2}+Q_{3}+Q_{4}=Mc\frac{\Delta T}{\Delta t}$, where $M$ is the mass and $c$ is the heat capacity of the CV, $\Delta T$ is the temperature change, and $\Delta t$ is the timestep interval.
@@
Figure 2: Thermal model for simulation. A building consists of conditioned zones, where the mean temperature of the zone $T_{z}$ should be within upper and lower setpoints, $\hat{T}_{z,max}$ and $\hat{T}_{z,min}$. Thermal power for heating or cooling the room is supplied to each zone, $\dot{Q}_{s}$, and recirculated from the zone, $\dot{Q}_{r}$ from the HVAC system, with additional thermal exchange $\dot{Q}_{z}$ from walls, doors, etc. The AHU supplies the building with air at supply air temperature setpoint $\hat{T}_{s}$ drawing fresh outside air, $\dot{m}_{OA}$, at temperatures, $T_{OA}$, and returning exhaust air $\dot{m}_{exhaust}$ at temperature $T_{exhaust}$ to the outside using intake and exhaust fans, $\dot{W}_{a,in}$ and $\dot{W}_{a,out}$. A fraction of the return air can be recirculated, $\dot{m}_{recirc}$. Central air conditioning is achieved with a chiller and pump that join a refrigeration cycle to the supply air, consuming electrical energy for the AC compressor $\dot{W}_{c}$ and coolant circulation, $\dot{W}_{c,p}$. The hot water cycle consists of a boiler that maintains the supply water temperature at $T_{b}$ heated by natural gas power $\dot{Q}_{b}$, and a pump that circulates hot water through the building, with electrical power $\dot{W}_{b,p}$. Supply air is delivered to zones through VAVs.
@@
The thermal exchange in (a) is calculated using Fourier’s law of steady conduction in the interior CVs (walls and interior air), parameterized by the conductivity of the CV, and the exchange across the exterior faces of CVs are calculated using the forced convection equation, parameterized by the convection coefficient, which approximates winds and currents surrounding the building. The change in internal energy (b) is parameterized by the density, and heat capacity of the CV. Finally, the thermal energy associated with the VAV (c) is equally distributed to all associated CVs that have a diffuser. Thermal diffusion within the building is mainly accomplished via forced or natural convection currents, which can be notoriously difficult to estimate accurately. We note that heat transfer using air circulation is effectively the exchange of air mass between CVs, which we approximate by a randomized shuffling of air within thermal zones, parameterized by a shuffle probability and radius. For further details see Appendix E.
@@
Simulator Configuration For RL to scale to many buildings, it is critical to be able to easily and rapidly configure the simulator to any arbitrary building. We designed a procedure that, given floorplans and HVAC layout information, enables generating a fully specified simulation very rapidly. For example, on SB1, consisting of 2 floors and 173 devices, a single technician was able to configure the simulator in under 3 hours. Details of this procedure are in Appendix F.
@@
Simulator Calibration and Evaluation In order to calibrate the simulator using sensor data, we must have a metric with which to evaluate our simulator’s fidelity, and an optimization method to improve our simulator on this metric.
@@
$\boldsymbol{N}$-Step Evaluation We propose a novel evaluation procedure, based on $N$-step prediction. Each iteration of our simulator was designed to represent a five-minute interval, and our real-world data is also obtained in five-minute intervals. To evaluate the simulator, we take a chunk of real data, consisting of $N$ consecutive observations. We then initialize the simulator so that its initial state matches that of the starting observation, and run the simulator for $N$ steps, replaying the same HVAC policy as was used in the real world. We then calculate our simulation fidelity metric, which is the mean absolute error of the temperatures in each temperature sensor at each timestep, averaged over time. More formally, we define the Temporal Spatial Mean Absolute Error (TS-MAE) of $Z$ zones over $N$ timesteps as:
@@
Where $T_{real,t,z}$ is measured zone air temperature for zone $z$ at timestamp $t$, and $T_{sim,t,z}=\frac{1}{|C_{z}|}\sum^{C_{z}}_{c=1}T_{t,c}$ is mean temperature of all CVs $C_{z}$ in zone $z$ at time $t$.
@@
Hyperparameter Calibration Once we define our simulation fidelity metric, TS-MAE, we can minimize this error, thus improving fidelity, by hyperparameter tuning several physical constants and other variables using black-box optimization methods. We chose the method outlined in Golovin et al. (2017), which automatically chooses the most appropriate strategy from a variety of popular algorithms.
5.1 Simulator Calibration Demonstration
We now provide an example of our calibration procedure.
Setup We configured the simulator to match SB1, with two stories, a combined surface area of 93,858 square feet, and 170 HVAC devices. Using the configuration pipeline, we went from floor plan blueprints to a fully configured simulator, a process that took a single technician less than three hours to complete. To calibrate, we took SB1 data from 3 days, from Monday July 10, 2023 12:00 AM, to Thursday July 13, 2023 12:00 AM. The first 2 days were used as a train set, and the third day as validation, as can be seen in Table 2. All times are local to the building.
Calibration Procedure We ran hyperparameter tuning for 4000 iterations to optimize the TS-MAE, as outlined in equation 1, on the training data. We reviewed the physical constants that yielded the lowest simulation error from calibration. Densities, heat capacities, and conductivities plausibly matched common interior and exterior building materials. However, the external convection coefficient is higher than under the weather conditions and likely is compensating for the radiative losses and gains, which were not directly simulated. For details about the hyperparameter tuning procedure, including the parameters varied, the ranges given, and the values found , see Appendix G.
Calibration Results In Table 2, we present the predictive results of our calibrated simulator, on $N$-step prediction, for the train scenario, where $N=576$, representing a two-day predictive window, and the test scenario, where $N=288$, representing a one day window. We calculated the TS-MAE, as defined in equation 1. We show results for the hyperparameters that best fit the train set, as well as for an uncalibrated simulator as a baseline. At no point was the validation data ever provided to the tuning process.
Table 2: Training and test data scenarios
Split Start End calib. $\epsilon$ uncalib. $\epsilon$
train 23-07-10 23-07-12 0.717 ${}^{\circ}C$ 1.971 ${}^{\circ}C$
val. 23-07-12 23-07-13 0.566 ${}^{\circ}C$ 1.618 ${}^{\circ}C$
As indicated in Table 2, our tuning procedure drifts only $0.56\circ C$ on average over a 24-hour validation period.
Visualizing Temperature Drift Over Time Figure 4 illustrates temperature drift over time for the training scenario. At each timestep, we calculate the spatial temperature for all sensors in both the real building and simulator and present them as side-by-side boxplot distributions for comparison. Figure 4 shows the same for the validation scenario.
@@ 5.1 Simulator Calibration Demonstration
Here we can see that our simulator temperature distribution maintains a minimal drift from the real world, although it does seem less reactive to daily fluctuation patterns, which may be due to the lack of a radiative heat transfer model.
@@
Figure 5: Visualization of simulator drift after 24 hours on validation data. A heat map represents the temperature difference between the simulator and the real world, with red indicating the simulator is hotter, blue indicating it is colder, and white indicating no difference. The zones with the max and min differences are indicated by displaying the difference above them.
Visualizing Spatial Errors Figure 5 illustrates the results of this predictive process over a 24-hour period, on the validation data. It displays a heatmap of the spatial temperature difference throughout the building, between the real world and simulator, after 24 hours of the simulator making predictions. The ring of blue around the building indicates that our simulator is too cold on the perimeter, which implies that the heat exchange with the outside is happening more rapidly than it would in the real world. The inside of the building, at least on the first floor, contains significant amounts of red, indicating that despite the simulator perimeter being cooler than the real world, the inside is warmer. This implies that our thermal exchange within the building is not as rapid as that of the real world. We suspect that this may be because our simulator does not have a radiative heat transfer model. Lastly, there is a large amount of white in this image, indicating that for the most part, even after 24 hours of making predictions on the validation data, our calibration process was successful and the fidelity remains high. For more visuals of spatial errors, see Appendix H.
6 The Smart Buildings PINN Model
Problem Formulation Recent work developed a modularized neural network model incorporating physical priors for building energy modeling, where different modules were designed to estimate distinct heat transfer terms in the dynamic building system.(Jiang & Dong, 2024). Building on this foundation, we updated the model structure to focus on control optimization. We consider the modeling task as a discrete-time dynamic system formulated in a state-space representation, with state variables, control inputs, and disturbance variables. As a control-oriented model, the primary focus is on managing model complexity while ensuring that its responses align with physical laws. To balance complexity and accuracy, we simplified the heat transfer terms by dividing them into three components: HVAC, adjacent zone, and other disturbances (e.g., outside air temperature, solar radiation, occupancy, and time features represented as sinusoidal functions). We further incorporated physical consistency constraints. The energy balance is expressed as:
$x(t+1)=x(t)+\frac{1}{cM}\Delta Q=x(t)+f_{NN_{A}}\Big{(}f_{NN_{B}}(u(t))\\ +f_{NN_{E}}(x(t),w(t))+f_{NN_{\text{adj}}}(x(t))\Big{)}$ (2)
Where $x$ represents the state variable (space air temperature), $M$ is the mass, and $c$ is the heat capacity of the space. $\Delta Q$ denotes the energy change within one timestep, which includes HVAC input and other heat transfer terms from conduction, convection, and radiation. $f_{NN_{A}}$, $f_{NN_{B}}$, $f_{NN_{E}}$, and $f_{NN_{\text{adj}}}$ are separate neural network modules to learn heat transfer dynamics, respectively.
Modular Design We develop an encoder-decoder structure for each neural network module. The encoder captures the thermal initial impact, providing a stable latent hidden state. Subsequently, we take a measurement of the current state to correct prediction errors at each timestep. The decoder then predicts outputs based on disturbance variables and control inputs. Beyond that, physical consistency is enforced through hard model parameter constraints to ensure that the control gain is positive. For example, additional cooling should decrease the space air temperature. This constraint, shown below, ensures that the partial derivative of the state variables to its control input is always positive before the current timestep. $\frac{\partial x_{t}}{\partial u_{k}}>0\quad\text{for}\quad k<t$
From Single-zone to Multi-zone To extend our model from a single-zone to a multi-zone framework, we integrated an adjacent heat transfer module to calculate conduction heat transfer based on temperature differences between zones. A skew-symmetric matrix, developed from the graph adjacency matrix, was used to represent this heat transfer term, ensuring that the heat transfer between two adjacent zones is equal in magnitude but opposite in direction.
Model Training The dataset used to train the model in these experiments was collected from SB4 between January 2023 and June 2024, and was split into training, validation, and testing sets with a ratio of 7:2:1. All features were normalized using a min-max scaler, mapping values to the range $[-1,1]$ except for the control input, which was scaled from $[0,1]$ due to physical consistency constraint concerns. To emulate a real-world training process, the size of the training data was gradually increased, starting with two days and extending to one and a half years, as more data would become available over time. To prevent overfitting, k-fold cross-validation and early stopping were employed.
Model Validation A well-predicting model does not necessarily ensure correct dynamic responses. To evaluate model performance, we used three metrics: 1) Mean Absolute Error (MAE): A standard metric to assess model accuracy. 2) Temperature Response Violation: At each timestep, we perform a sanity check by applying additional heating or cooling to verify if the direction of temperature change is correct. If the sign is correct, the violation is zero. Otherwise, we calculate and accumulate the temperature violation over time. 3) Maximum mean discrepancy (MMD): A distance metric on the space of probability measures. MMD quantifies the difference between two sets of samples by taking the maximum difference in sample averages over a kernel function. We use MMD to evaluate the similarity between the model’s responses and the ground truth responses collected from measured data.
Figure 6: ModNN model performance on an example day.
Model Performance Figure 6 illustrates the model performance on an example day. The gray line represents the predicted space air temperature, which closely matches the measured data, achieving an MAE of 0.25 °C for 24-hour predictions. The red and blue lines show the sanity check results, with red indicating maximum heating and blue indicating maximum cooling. The model’s responses align with underlying physical laws, and the temperature response violation is zero. This demonstrates that the physics-informed constraints effectively ensure accurate and physically consistent model responses. We also compared the MMD of the LSTM and ModNN models, as illustrated in Figure 7. The X-axis represents one-step HVAC load changes (negative values indicate increased cooling, positive values indicate reduced cooling), and the Y-axis shows the corresponding changes in space air temperature. Each scatter point reflects data under varying weather and occupancy conditions.
Figure 7: Maximum mean discrepancy of LSTM and ModNN.
From the black points (ground truth), we observe a clear decreasing trend in space air temperature as the cooling load increases. The ModNN model, represented by blue points, captures a similar trend. However, the LSTM model, depicted by red points, shows an incorrect trend where the space air temperature increases with additional cooling. This discrepancy is further highlighted by the difference in distributions between the ground truth (black contour plot) and the LSTM model (red contour plot). The MMD of the LSTM model to the ground truth is 0.14, which is significantly higher than the MMD of the ModNN model to the ground truth (0.05). For more details, see Appendix I.
7 Learning an Improved Control Policy